OpenStudy (anonymous):

Find the quotient: 49a^2-b^2 / 7a-b

OpenStudy (anonymous):

\[\frac{49 a^2-b^2}{7 a-b}=\frac{(7 a-b) (7 a+b)}{7 a-b}= (7 a+b) \]

OpenStudy (anonymous):

thank you, I was off just a slight bit.. will go back and review

OpenStudy (anonymous):

You are welcome.